OpenStudy (anonymous):

racing car traveling with constant acceleration increases its speed from 10 m/s to 50 m/s over a distance of 60 m. How long does this take?

OpenStudy (irishboy123):

use: \(v^2 = u^2 + 2 a x\) to get 'a' then: \(v = u + at\) to get t

OpenStudy (anonymous):

I had also done exactly like that but this can't give the appropriate answer

OpenStudy (irishboy123):

from (1) \(a= \frac{50^2 - 10^2}{2 \times 60} = 20m/s^2\) and (2) \(t = \frac{50 - 10}{20} = 2 s\) to be sure, check using final equation of motion: \(x = ut + \frac{1}{2}a t^2 = 10(2) + \frac{1}{2} (20) 2^2 = 60m \ \huge \checkmark\)

OpenStudy (anonymous):

Hmmm thanks brother...me too got the same answer but the book from which i isolated this question have 4s answer in it thats why i was just confirming it